4. Light, with 550 nm wavelength, is used to illuminate double slits which are separated of 3.5 x 10-5 m. A screen is placed 1.25 m from the slits.

   a. Find the angle of the first bright fringe.

   b. Find the angle of the third bright fringe.
Answer #1

a)

as we know

d sin x = nL

3.5* 10^-5 sin x = 1* 550* 10^-9

x = 0.9 deg

b)

similarly

3.5* 10^-5 sin x = 3* 550* 10^-9

x = 2.7 deg

======

Comment before rate in case any doubt, will reply for sure.. goodluck